# Positive formula for the product of conjugacy classes on the unitary group

### v1

@@ 6.5. Quasi dual hives
For two color maps $C$, $C^{\prime}$ that differ by more than one hexagon rotation, we denote $Rot[C\rightarrow C^{\prime}]$ the composition of maps in 6.19 for each hexagon rotation needed to go from $C$ to $C_{0}$ and then from $C_{0}$ to $C^{\prime}$ where is existence of such paths was given by Proposition 6.14.
@@
Let $H=(C_{0},L)\in\widetilde{H}^{C_{0}}(\lambda,\mu,\nu,N)$ be a quasi hive having the simple color map $C_{0}$. Let $I\subset E_{n,d}$ be the set of edges of $E_{n,d}$ that are not in a rigid lozenge of $C_{0}$ except the edges of type $2$ between a rigid lozenge and a triangular face with colors $(0,0,0)$ and edges of type $1$ between a rigid lozenge and a triangular face with colors $(1,1,1)$, see Figure 30 below for an example. Denote $I_{0}$ the edges of $I$ of type 00 where we remove the east-most such edge on each row. There are $D$ such edges so that $I_{0}=(e_{1},\dots,e_{D})$.
Then, for any $z=(z_{1},\dots,z_{D})\in\left(\frac{1}{N}\mathbb{Z}\right)^{D}$, there exists a unique label map $\Phi^{C_{0}}(z)\in L^{C_{0}}(\lambda,\mu,\nu,N)$ such that for all $1\leq i\leq D$: $\Phi^{C_{0}}(z)(e_{i})=z_{i}$. Moreover, for all $1\leq i\leq D$, $\Phi^{C_{0}}(z)(e_{i})$ is given by an affine combination with integer coefficients of $(z_{1},\dots,z_{D},\lambda,\mu,\nu,\frac{1}{N})$.
@@
Let $L:\partial E_{n,d}\rightarrow\frac{1}{N}\mathbb{Z}$ be a function defined on the boundaries of $E_{n,d}$ satisfying the boundary condition $(\lambda,\mu,\nu,N)$. We will show that $L$ can be extend to a quasi label $E_{n,d}$ in a unique way. For any edge of type $2$ part of an rigid lozenge, there exists an edge $e_{\partial}(e)\in\partial E_{n,d}$ of the same type obtained by translation of $e$ by a multiple of $\mathrm{e}^{i\frac{2\pi}{3}}$. Likewise, for any edge of type $1$ part of an rigid lozenge, there exists an edge $e_{\partial}(e)\in\partial E_{n,d}$ of the same type obtained by translation of $e$ by a multiple of $\mathrm{e}^{i\frac{\pi}{3}}$. Assign $L(e)=L(e_{\partial}(e))$ for each such edge $e$. By the equality condition on opposite edges of rigid lozenges, any quasi label map has the same values on these edges.
It remains to extend $L$ to edges $e\in I$. The values of $\partial I$ are already determined uniquely by the boundary conditions. Set $L(e_{i})=z_{i}$ for $1\leq i\leq D$. We call a band the following configuration of adjacent faces where the west-most face has both its edges of type 00 and $2$ labeled, the east-most face has its type $1$ edge labeled and all other faces in between have their type 00 edge labeled.
Figure 31. A band of size four. Marked edges are the already labeled edge. The edge labels of other edges are determined in the order of the red numbers by face summation constraints.
@@
The previous paragraph show that one can extend $L$ to all edges in the south most band of $I$. Notice then that the region above is also a band. By induction, one extends $L$ to the $n-d$ south most bands of $I$. The remaining region consists of bands turned upside down which can are also uniquely determined by the same reasoning. The resulting map $L$ on $E_{n,d}$ satisfies the face summation constraints so that $L\in\widetilde{L}^{C_{0}}(\lambda,\mu,\nu,N)$. Since the band completions are unique at each step, the labels of any other quasi label map $L^{\prime}\in\widetilde{L}^{C_{0}}(\lambda,\mu,\nu,N)$ with same values on $I_{0}$ would agree with $L$. ∎
@@
Define $\Phi^{C_{0}}(z)\in L^{C_{0}}(\lambda,\mu,\nu)$ to be the unique label map constructed in Lemma 6.21 from specifying edge coordinates $z\in\left(\frac{1}{N}\mathbb{Z}\right)^{D}$. Let $C:E_{n,d}\rightarrow\{0,1,3,m\}$ be any color map and let $D=(n-1)(n-2)/2$. Define the following map
$\Phi^{C}:\left(\frac{1}{N}\mathbb{Z}\right)^{D}$ $\rightarrow L^{C}(\lambda,\mu,\nu,N)$
@@
The map $\Phi^{C}$ of Definition 6.22 is bijective. Moreover, for every $z\in\left(\frac{1}{N}\mathbb{Z}\right)^{D}$ and edge $e\in E_{n,d}$ , $\Phi^{C}(z)(e)$ is an affine combination of $(z_{1},\dots,z_{D},\lambda,\mu,\nu,\frac{1}{N})$ with integer coefficients.
@@
Let $C,C^{\prime}$ be two color maps. We extend the maps $Rot[C\rightarrow C^{\prime}]$ and $\Phi^{C}$ of Definitions 6.19 and 6.22 to quasi hives by
@@ 7.1. Limit dual hives
      $L$ satisfies the same inequalities as label maps for dual hives stated in $(2)$ of Definition 5.2 replacing the strict inequalities in $(2c)$ by a large one.
@@
      The values of $L$ on $\partial E_{n,d}$ are given by $(\alpha,\beta,\gamma)$.
As in the discrete case, if $C$ is a color map, we denote by $H^{C}(\lambda,\mu,\nu,\infty)\subset H(\lambda,\mu,\nu,\infty)$ the subset of limit dual hives of having color map $C$. As in the previous section, we denote by $\widetilde{H}(\lambda,\mu,\nu,\infty)$ the set of pairs $(C,L)$ as above where we remove the inequality conditions $(2b)$ on $L$.
@@
Note that Lemmas 6.18, 6.20 and 6.21 hold for limit quasi dual hives $H=(C,L)\in\widetilde{H}(\lambda,\mu,\nu,\infty)$ extending $z\in\left(\frac{1}{N}\mathbb{Z}\right)^{D}$ to $z\in\mathbb{R}^{D}$. Using the same construction as in Section 6, we define $Rot[C\rightarrow C^{\prime}]:\widetilde{H}^{C}(\lambda,\mu,\nu,\infty)\rightarrow\widetilde{H}^{C^{\prime}}(\lambda,\mu,\nu,\infty)$ and $\Phi^{C}:\mathbb{R}^{D}\rightarrow\widetilde{H}^{C}(\lambda,\mu,\nu,\infty)$ which are affine bijections with integer coefficients.
@@ 7.2. Convergence to a volume
The goal of this part is to prove Proposition 7.3 which expresses the limit quantum cohomology coefficients as the volume involving limit dual hives. The proof relies on Lemma 7.4 Lemma 7.6 below.
@@
Recall that
@@
(7.3) $=\sum_{C}\int_{\mathbb{R}^{D}}\sum_{z=(z_{1},\dots,z_{d})\in(\frac{1}{N}\mathbb{Z})^{D}:\Phi^{C}(z)\in\widetilde{H}(\lambda_{N},\mu_{N},\nu_{N},N)\cap Ineq(n)}\mathbf{1}(u)_{z+[-\frac{1}{N},\frac{1}{N}]^{D}}du.$
@@
Then, for any $u\in\mathbb{R}^{D}$:
(7.4) $\lim_{N\rightarrow\infty}f_{N}^{C}(u)=\mathbf{1}(u)_{\{\Phi^{C}[u]\in H^{C}(\lambda,\mu,\nu,\infty)\}}.$
@@
Note that a priori, $\Phi^{C}[u]\in\widetilde{H}^{C}(\lambda,\mu,\nu,\infty)$ is a label map such that $(C,L)$ is a limit hive of Definition 7.1 without inequality constraints $(2c)$ and (2d). Here, the right hand side is more restrictive as in requires that $(C,\Phi^{C}[u])\in H^{C}(\lambda,\mu,\nu,\infty)=\widetilde{H}^{C}(\lambda,\mu,\nu,\infty)\cap Ineq(n)$.
@@
Take $u$ such that $\Phi^{C}[u]$ in the interior of $H^{C}(\lambda,\mu,\nu,\infty)$. We want to show that $f_{N}^{C}(u)=1$ for $N\geq N_{0}$ which means that one can find a sequence $(z^{(N)},N\geq N_{0})=((z_{1}^{(N)},\dots,z_{D}^{(N)}),N\geq N_{0})$ such that for each $N$ : $u\in z_{N}\pm 1/N$. Let us take the label map
@@
associated to $z^{(N)}=\lfloor Nu\rfloor/N$ : $L^{(N)}=\Phi^{C}_{N}(z^{(N)})$ where $\Phi^{C}_{N}$ is associated to the boundaries $(\lambda_{N},\mu_{N},\nu_{N})$. One has $|z^{(N)}-u|\leq 1/N$ by construction. We need to check that $L^{(N)}\in\widetilde{H}(\lambda_{N},\mu_{N},\nu_{N},N)\cap Ineq(n)$.
By definition, $\Phi^{C}_{N}$ is a quasi label map with boundary conditions $(\lambda_{N},\mu_{N},\nu_{N})$ so that $L^{(N)}\in\widetilde{H}(\lambda_{N},\mu_{N},\nu_{N},N)$. Let us check the inequality constraints of $Ineq(n)$. Take any pair of edges $(e,e^{\prime})$ subject to an inequality. Since $\Phi^{C}[u]\in H^{C}(\lambda,\mu,\nu,\infty)$ is in the interior, this equality is sharp for $\Phi^{C}[u]$ that is
@@
Taking $N_{0}=\max(N_{0}(e,e^{\prime}),(e,e^{\prime})\in E_{n,d}^{2})$, $L^{(N)}$ satisfies the inequality constraints for $N\geq N_{0}$. Therefore,
@@
so that $\lim_{N\rightarrow+\infty}f_{N}^{C}(u)=1$ as desired. For $\Phi^{C}[u]\notin H^{C}(\lambda,\mu,\nu,\infty)$, at least one of the two conditions $(2c),(2d)$ of Definition 7.1 is violated. It cannot be that conditions $(2a),(2b)$ do not hold by construction of $\Phi^{C}$. Let $(e,e^{\prime})$ be a pair of edge such that either $(2c)$ or $(2d)$ is violated : $\Phi^{C}[u](e)<\Phi^{C}[u](e^{\prime})$. Using that $\lim_{N\rightarrow+\infty}L^{(N)}(e)=\Phi^{C}[u](e)$, one has that for $N$ large enough $L^{(N)}(e)<L^{(N)}(e^{\prime})$ so that $L^{(N)}\notin\widetilde{H}(\lambda_{N},\mu_{N},\nu_{N},N)\cap Ineq(n)$. Therefore, we have that (7.4) holds.
@@
Let $f_{N}^{C}(u)$ be as in (7.4). Then, there exists a compact $K^{C}\subset\mathbb{R}^{D}$ such that for every $N\geq 1$,
(7.9) $|f_{N}^{C}(u)|\leq\mathbf{1}_{\{u\in K^{C}\}}.$
@@
If $C=C_{0}$, the values $z\in\mathbb{Z}^{D}$ such that $\Phi_{N}^{C_{0}}(z)\in\widetilde{H}^{C_{0}}(\lambda_{N},\mu_{N},\nu_{N},N)\cap Ineq(n)$ are in $[0,1]^{D}$ since $(z_{1},\dots,z_{D})$ are the values of $(\Phi_{N}^{C_{0}}(z)(e_{1}),\dots,\Phi_{N}^{C_{0}}(z)(e_{D}))$ for some horizontal edges $(e_{1},\dots,e_{D})\in E_{n,d}$ which are in $[0,1]$.
If $C\neq C_{0}$, then there exists a composition of rotation $Rot[C\rightarrow C_{0}]$ such that $Rot[C\rightarrow C_{0}](e)\in[0,1]$ for each horizontal edge $e$ so that $z\in\cup_{e}\left(z\mapsto Rot[C\rightarrow C_{0}](e)^{-1}[0,1]\right)$ which is compact.
@@
where $\Phi^{C_{0}}(z)$ is the label map of the quasi hive with simple color map $C_{0}$ having horizontal edge labels $z$. For $z\in\mathbb{Z}^{D}$ such that $\Phi_{N}^{C}(z)\in\widetilde{H}^{C}(\lambda_{N},\mu_{N},\nu_{N},N)\cap Ineq(n)$, we know from $Ineq(n)$ that values $\{\Phi_{N}^{C}(z)(e),e\in E_{n,d}\}$ are in $[0,1]^{E_{n,d}}$. Applying the affine hence continuous $Rot[C\rightarrow C_{0}]$ we get that $Rot[C\rightarrow C_{0}]\Phi^{C}(z)=\Phi^{C_{0}}(z)\in Rot[C\rightarrow C_{0}]([0,1]^{E_{n,d}})$ which is compact. In particular, the labels $(z_{i}=\Phi^{C_{0}}(z)(e_{i}),1\leq i\leq D)$ of horizontal edges $e_{i}$ in $I$ are in compact sets hence bounded.
@@
By Lemma 7.4 Lemma 7.6, using dominated convergence in (7.3):

### v2

@@ 6.5. Quasi dual hives
For two color maps $C$, $C^{\prime}$ that differ by more than one hexagon rotation, we denote $Rot[C\rightarrow C^{\prime}]$ the composition of maps in Definition 6.22 for each hexagon rotation needed to go from $C$ to $C_{0}$ and then from $C_{0}$ to $C^{\prime}$ where the existence of such paths was given by Proposition 6.16. Note that there might be multiple rotation paths from $C$ to $C^{\prime}$ so that such a map is not unique.
Let $I\subset E_{n,d}$ be the set of edges of $E_{n,d}$ that are not in a rigid lozenge of $C_{0}$ except the edges of type $2$ between a rigid lozenge and a triangular face with colors $(0,0,0)$ and edges of type $1$ between a rigid lozenge and a triangular face with colors $(1,1,1)$, see Figure 25 below for an example. Denote $I_{0}$ the edges of $I$ of type 00 where we remove the east-most such edge on each row. By a counting argument, there are $D=\frac{(n-1)(n-2)}{2}$ such edges so that $I_{0}=(e_{1},\dots,e_{D})$.
@@
For any $z=(z_{1},\dots,z_{D})\in\left(\frac{1}{N}\mathbb{Z}\right)^{D}$, there exists a unique label map $\Phi^{C_{0}}(z)\in L^{C_{0}}(\lambda,\mu,\nu,N)$ such that for all $1\leq i\leq D$: $\Phi^{C_{0}}(z)(e_{i})=z_{i}$. Moreover, for all $1\leq i\leq D$, $\Phi^{C_{0}}(z)(e_{i})$ is given by an affine combination with integer coefficients of $(z_{1},\dots,z_{D},\lambda,\mu,\nu,\frac{1}{N})$.
@@
Let $L:\partial E_{n,d}\cup I_{0}\rightarrow\frac{1}{N}\mathbb{Z}$ be a function satisfying the boundary condition $(\lambda,\mu,\nu,N)$ as in (6.7). We will show that $L$ can be extended to a quasi label $E_{n,d}$ in a unique way. For any edge of type $2$ part of an rigid lozenge, there exists an edge $e_{\partial}(e)\in\partial E_{n,d}$ of the same type obtained by translation of $e$ by a multiple of $\mathrm{e}^{i\frac{2\pi}{3}}$. Likewise, for any edge of type $1$ part of an rigid lozenge, there exists an edge $e_{\partial}(e)\in\partial E_{n,d}$ of the same type obtained by translation of $e$ by a multiple of $\mathrm{e}^{i\frac{\pi}{3}}$. Assign $L(e)=L(e_{\partial}(e))$ for each such edge $e$. By the equality condition on opposite edges of rigid lozenges, any quasi label map has the same values on these edges.
It remains to extend $L$ to edges $e\in I$. The values of $\partial I$ are already determined uniquely by the boundary conditions. Set $L(e_{i})=z_{i}$ for $1\leq i\leq D$. We call a band the following configuration of adjacent faces where the west-most triangular face has both its edges of type 00 and $2$ labeled, the east-most triangular face has its type $1$ edge labeled and all other faces in between have their type 00 edge labeled.
Figure 26. A band of size four. Marked edges are the already labeled edges. The labels of other edges are determined in the order of the red numbers by face summation constraints.
@@
The previous paragraph shows that one can extend $L$ to all edges in the south most band of $I$. Notice then that the region above is also a band. By induction, one extends $L$ to the $n-d$ south most bands of $I$. The remaining region consists of bands turned upside down which are also uniquely determined by the same reasoning. The resulting map $L$ on $E_{n,d}$ satisfies the face summation constraints so that $L\in\widetilde{L}^{C_{0}}(\lambda,\mu,\nu,N)$. Since the band completions are unique at each step, the labels of any other quasi label map $L^{\prime}\in\widetilde{L}^{C_{0}}(\lambda,\mu,\nu,N)$ with same values on $I_{0}$ would agree with $L$. ∎
@@
Define $\Phi^{C_{0}}(z)\in L^{C_{0}}(\lambda,\mu,\nu)$ to be the unique label map constructed in Lemma 6.24 from specifying edge coordinates $z\in\left(\frac{1}{N}\mathbb{Z}\right)^{D}$ on $I_{0}$. Let $C:E_{n,d}\rightarrow\{0,1,3,m\}$ be any color map and let $D=(n-1)(n-2)/2$. Define the following map
$\Phi^{C}:\hskip 42.67912pt\left(\frac{1}{N}\mathbb{Z}\right)^{D}$ $\rightarrow L^{C}(\lambda,\mu,\nu,N)$
@@
The map $\Phi^{C}$ of Definition 6.25 is bijective. Moreover, for any $z\in\left(\frac{1}{N}\mathbb{Z}\right)^{D}$ and edge $e\in E_{n,d}$ , $\Phi^{C}(z)(e)$ is an affine combination of $(z_{1},\dots,z_{D},\lambda,\mu,\nu,\frac{1}{N})$ with integer coefficients.
@@
Let $C,C^{\prime}$ be two color maps. We extend the maps $Rot[C\rightarrow C^{\prime}]$ and $\Phi^{C}$ of Definitions 6.22 and 6.25 to quasi hives by
@@ 7. Convergence to a volume of hives
Since $\mathrm{d}\mathbb{P}[\gamma+2r|\alpha+r,\beta+r]=\mathrm{d}\mathbb{P}[\gamma|\alpha,\beta]$ for $r>0$, assume in this section without loss of generality that $\alpha,\beta,\gamma\in\mathcal{H}_{reg}$ are such that $\alpha_{n},\beta_{n}>0$ and $\gamma_{1}<1$. In particular, if $\lambda_{N},\mu_{N},\nu_{N}$ are such that $\frac{1}{N}\lambda_{N}\rightarrow\alpha,\frac{1}{N}\mu_{N}\rightarrow\beta,\frac{1}{N}\nu_{N}\rightarrow\gamma$, then $\min((\lambda_{N})_{n},(\mu_{N})_{n},N-n-(\nu_{N})_{1})>d+1$ for $N$ large enough. Hence, we will assume throughout this section that the hives with boundary $(\lambda_{N},\mu_{N},\nu_{N})$ are regular in the sense of (6.7), see Figure 22.
@@ 7.1. Limit dual hives
This section introduces limit dual hives which will be linked to the limit of quantum Littlewood-Richardson coefficients of Theorem 3.8.
@@
      if $e,e^{\prime}$ are edges of same type on the boundary of a same lozenge $f$,
      1. (i)
        $L(e)=L(e^{\prime})$ if the middle edge of $f$ is colored $m$,
      2. (ii)
        $L(e)\geq L(e^{\prime})$ if $h(e)>h(e^{\prime})$.
@@
      The values of $L$ on $\partial E_{n,d}$ are given by $(\alpha,\beta,\gamma)$ so that, sorted in decreasing height of edges, see Figure 27 below.
      $\ell^{(0,1)}$ $=(1-\alpha_{d},\dots,1-\alpha_{1}),$ $\ell^{(2,2)}$ $=(\alpha_{d+1},\dots,\alpha_{n})$
      $\ell^{(2,0)}$ $=(1-\beta_{d},\dots,1-\beta_{1}),$ $\ell^{(1,1)}$ $=(\beta_{n},\dots,\beta_{d+1})$
      $\ell^{(1,2)}$ $=(\gamma_{n},\dots,\gamma_{n-d+1}),$ $\ell^{(2,2)}$ $=(1-\gamma_{n-d},\dots,1-\gamma_{1}).$
Figure 27. Boundary labels for limit hives in $H(\alpha,\beta,\gamma,\infty)$.
As in the discrete case, if $C$ is a color map, we denote by $H^{C}(\lambda,\mu,\nu,\infty)\subset H(\lambda,\mu,\nu,\infty)$ the subset of limit dual hives of having color map $C$. As in the previous section, we denote by $\widetilde{H}(\lambda,\mu,\nu,\infty)$ the set of pairs $(C,L)$ as above where we remove the inequality conditions $(2.b.ii)$ on $L$.
@@
Note that Lemmas 6.21, 6.23 and 6.24 hold for limit quasi dual hives $H=(C,L)\in\widetilde{H}(\lambda,\mu,\nu,\infty)$ extending $z\in\left(\frac{1}{N}\mathbb{Z}\right)^{D}$ to $z\in\mathbb{R}^{D}$. Using the same construction as in Section 6, we define $Rot[C\rightarrow C^{\prime}]:\widetilde{H}^{C}(\lambda,\mu,\nu,\infty)\rightarrow\widetilde{H}^{C^{\prime}}(\lambda,\mu,\nu,\infty)$ and $\Phi^{C}:\mathbb{R}^{D}\rightarrow\widetilde{H}^{C}(\lambda,\mu,\nu,\infty)$ which are affine bijections with coefficients in $\mathbb{Z}[\alpha,\beta,\gamma]$ obtained by setting $\frac{1}{N}=0$.
@@ 7.2. Convergence to a volume
The goal of this part is to prove Proposition 7.3 which expresses the limit quantum cohomology coefficients as the volume involving limit dual hives. The proof relies on Lemma 7.4 and Lemma 7.6 below.
@@
Recall from Corollary 5.4 that
@@
(7.3) $=\sum_{C}\int_{\mathbb{R}^{D}}\sum_{z=(z_{1},\dots,z_{d})\in(\frac{1}{N}\mathbb{Z})^{D}:\Phi^{C}(z)\in\widetilde{H}(\lambda_{N},\mu_{N},\nu_{N},N)\cap Ineq(n)}\mathbf{1}(u)_{z+[-\frac{1}{N},\frac{1}{N}[^{D}}du.$
@@
Recall that $\frac{1}{N}\lambda_{N}=\alpha+o(1)$, $\frac{1}{N}\mu_{N}=\beta+o(1)$ and $\frac{1}{N}\nu_{N}=\gamma+o(1)$ as $N\rightarrow+\infty$. Then, for almost all $u\in\mathbb{R}^{D}$ with respect to the Lebesgue measure:
(7.4) $\lim_{N\rightarrow\infty}f_{N}^{C}(u)=\mathbf{1}(u)_{\{\Phi^{C}[u]\in H^{C}(\alpha,\beta,\gamma,\infty)\}}.$
@@
Note that a priori, $\Phi^{C}[u]\in\widetilde{H}^{C}(\alpha,\beta,\gamma,\infty)$ is a label map such that $(C,L)$ is a limit hive of Definition 7.1 without inequality constraints $(2c)$ and (2d). Here, the right hand side is more restrictive as it requires that $(C,\Phi^{C}[u])\in H^{C}(\alpha,\beta,\gamma,\infty)=\widetilde{H}^{C}(\alpha,\beta,\gamma,\infty)\cap Ineq(n)$.
@@
Take $u$ such that $\Phi^{C}[u]$ in the interior of $H^{C}(\alpha,\beta,\gamma,\infty)$. We want to show that $f_{N}^{C}(u)=1$ for $N\geq N_{0}$ which means that one can find a sequence $(z^{(N)},N\geq N_{0})=((z_{1}^{(N)},\dots,z_{D}^{(N)}),N\geq N_{0})$ such that for each $N\geq N_{0}$ : $\Phi^{C}(z^{(N)})\in\widetilde{H}(\lambda_{N},\mu_{N},\nu_{N},N)\cap Ineq(n)$ and $u\in z^{(N)}+[-1/N,1/N[$. Let us take the label map
@@
associated to $z^{(N)}=\lfloor Nu\rfloor/N$ : $L^{(N)}=\Phi^{C}_{N}(z^{(N)})$ where $\Phi^{C}_{N}$ is associated to the boundaries $(\lambda_{N},\mu_{N},\nu_{N})$. One has $|z^{(N)}-u|<1/N$ by construction. We need to check that $L^{(N)}\in\widetilde{H}(\lambda_{N},\mu_{N},\nu_{N},N)\cap Ineq(n)$.
By definition, $\Phi^{C}_{N}$ is a quasi label map with boundary conditions $(\lambda_{N},\mu_{N},\nu_{N})$ so that $L^{(N)}\in\widetilde{H}(\lambda_{N},\mu_{N},\nu_{N},N)$. Let us check the inequality constraints of $Ineq(n)$. Take any pair of edges $(e,e^{\prime})$ subject to an inequality. Since $\Phi^{C}[u]\in H^{C}(\alpha,\beta,\gamma,\infty)$ is in the interior, this equality is sharp for $\Phi^{C}[u]$ that is
@@
Hence, $L^{(N)}$ satisfies all the inequality constraints for $N\geq N_{0}$, where $N_{0}$ is the largest of the thresholds $N_{0}(e,e^{\prime})$ for $(e,e^{\prime})$ related by an inequality constraint in Definition 7.1. Therefore,
@@
so that $\lim_{N\rightarrow+\infty}f_{N}^{C}(u)=1$ as desired. For $\Phi^{C}[u]\notin H^{C}(\alpha,\beta,\gamma,\infty)$, one of the inequalities in $(2.b.ii)$ of Definition 7.1 is violated, for all other conditions being satisfied by construction of $\Phi^{C}$. Let $(e,e^{\prime})$ be a pair of edge such that $(2.b.ii)$ is not satisfied : $\Phi^{C}[u](e)<\Phi^{C}[u](e^{\prime})$ while $h(e)>h(e^{\prime})$ for some pair of edges of same type adjacent to a same lozenge. Using that $\lim_{N\rightarrow+\infty}L^{(N)}(e)=\Phi^{C}[u](e)$, one has that for $N$ large enough $L^{(N)}(e)<L^{(N)}(e^{\prime})$ so that $L^{(N)}\notin\widetilde{H}(\lambda_{N},\mu_{N},\nu_{N},N)\cap Ineq(n)$. Hence, (7.4) holds for almost all $u\in\mathbb{R}^{D}$ with respect to the Lebesgue measure.
@@
Let $f_{N}^{C}(u)$ be as in (7.4). Then, there exists a compact $K^{(C)}\subset\mathbb{R}^{D}$ such that for every $N\geq 1$,
(7.9) $|f_{N}^{C}(u)|\leq\mathbf{1}_{\{u\in K^{(C)}\}}.$
@@
If $C=C_{0}$, the values $z\in\mathbb{Z}^{D}$ such that $\Phi_{N}^{C_{0}}(z)\in\widetilde{H}^{C_{0}}(\lambda_{N},\mu_{N},\nu_{N},N)\cap Ineq(n)$ are in $[0,1]^{D}$ since $(z_{1},\dots,z_{D})$ are the values of $(\Phi_{N}^{C_{0}}(z)(e_{1}),\dots,\Phi_{N}^{C_{0}}(z)(e_{D}))$ for some horizontal edges $(e_{1},\dots,e_{D})\in E_{n,d}$ which are in $[0,1]$ by construction.
@@
where $\Phi^{C_{0}}(z)$ is the label map of the quasi hive with simple color map $C_{0}$ having horizontal edge labeled $z$. For $z\in\mathbb{Z}^{D}$ such that $\Phi_{N}^{C}(z)\in\widetilde{H}^{C}(\lambda_{N},\mu_{N},\nu_{N},N)\cap Ineq(n)$, we know from $Ineq(n)$ that values $\{\Phi_{N}^{C}(z)(e),e\in E_{n,d}\}$ are in $[0,1]^{E_{n,d}}$. Applying the affine hence continuous map $Rot[C\rightarrow C_{0}]$, we get that $Rot[C\rightarrow C_{0}]\Phi^{C}(z)=\Phi^{C_{0}}(z)\in Rot[C\rightarrow C_{0}]([0,1]^{E_{n,d}})$ which is compact. In particular, the labels $(z_{i}=\Phi^{C_{0}}(z)(e_{i}),1\leq i\leq D)$ of horizontal edges $e_{i}$ in $I$ are in compact sets hence bounded.
@@
By Lemma 7.4 and Lemma 7.6, using dominated convergence theorem in (7.3):